5 mangoes are placed in a fruit container in Mrs Jeremiah’s home. In the fruit container, 3 are green, one is red and the last is yellow.

a) What is the probability that 3 are pulled one after the order without replacement and they are in order of green mango, not a green mango, green mango?

b) A mango is pulled, and it is either a red mango or a yellow mango

c) Two mangoes are pulled one after the other without replacement and none of them is a green mango

d) What is the probability that 3 are pulled one after the order replacement and they are in order of green mango, not a green mango, green mango?
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

a)

Order - Green, Red/Yellow, Green

=> Prob (green mango, not a green mango, green mango) = 3C1*2C1*2C1/ 5*4*3 = 0.2

b)

Prob (red mango or a yellow mango) = 2C1/5C1 = 0.4

c)

Prob (no green) = 2C1*1C1/5*4 = 0.1

d)

Prob(green mango, not a green mango, green mango) = 3C1*2C1*3C1/5*5*5 = 0.144